OpenStudy (anonymous):

what is the derivative of -4/x^5+(8*3sqrtx)+8

OpenStudy (amistre64):

this is just a simple application of the derivative rules. what have you attempted so far?

OpenStudy (anonymous):

I have tried to bring the x^5 up next to the -4 so it would read -4x^-5, but that did not work. i tried to find the lcd but that didnt work either as well.

OpenStudy (amistre64):

some things to keep in mind:\[\frac{d}{dx}x^v=n*x^{n-1}\] \[\sqrt{a}=a^{1/2}\] \[\frac{1}{b}=b^{-1}\]

OpenStudy (amistre64):

what is your definition of "did not work"?

OpenStudy (anonymous):

sorry. I have to input what i think is the answer on an online database. When i typed in what i thought the answer was, it was wrong.

OpenStudy (amistre64):

oh, yeah online programs are the worst for getting a proper format ....

OpenStudy (amistre64):

-4/x^5+(8*3sqrtx)+8 -4x^-5 +8*3x^1/2 + 8 -4.-5x^(-5-1) + 8.3.1/2 x^(1/5-1) + 8.0

OpenStudy (amistre64):

the . is a lazy *

OpenStudy (amistre64):

1/2 - 1 .... not 1/5 - 1 as well :)

OpenStudy (anonymous):

thank you, i will try that and see if it works. thank you

OpenStudy (mertsj):

Did it work?